bianca is twice as old as pio and margie as twice as old as bianca. In 10 years, their combined ages is 58.How old is bianca now?
Answer by CubeyThePenguin(3113)

B = Bianca's age
P = Pio's age
M = Margie's age

B = 2P
M = 2B
(B+10)+(P+10)+(M+10) = 58

Write everything in terms of B.

B + P + M = 28
2P + P + 4P = 28
P = 4

Bianca is twice as old as Pio, so Bianca is 8.